OpenStudy (anonymous):

HELP I need help with a couple algebra problems? These are the ones I have gotten wrong so far and I just cant figure them out to save my life. I have this class online so I get the chance to redo them for full credit. 13.) 3x-7/10 ≥ 1/3x+7 8.) x= -8/9y+5/8 ; for y 10.) solve for X 5(6ax+y) = 29ax-3y 12.) |2.3-0.4x|=1 13.) |1-3x/7| = 2/3 21.) 7x+9 ≤ 3x - 11 14.) |8x-7/9| = |8x+2| 15.) |3-x| = |x/2+p|

OpenStudy (anonymous):

13) First subtract 1/3x from both sides, then add 7/10 to both sides, then divide both sides divide both sides by 2 2/3, or 8/3. Is that what you did?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And you got that wrong? :o

OpenStudy (anonymous):

What did you get when that was all said and done? I got \[x ≥ 2 \frac{71}{80}\]

OpenStudy (anonymous):

it wont let me see the wrong answer i got but i know i did not get that...

OpenStudy (anonymous):

my class said that answer is wrong too...

OpenStudy (anonymous):

@NinjaDevo

OpenStudy (texaschic101):

13. 3x - 7 /10 >= 1/3x + 7 --- multiply everything by common denominator (30) to get rid of fractions 90x - 21 >= 10x + 210 --- subtract 10x from both sides 90x - 10x - 21 >= 210 --- add 21 to both sides 90x - 10x >= 210 + 21 --- combine like terms 80x >= 231 -- divide both sides by 80 x >= 231/80 or x = 2 71/80 thats what I got :) =============== 8. x = -8/9y + 5/8 (like I said above, if you want to get rid of fractions, multiply by the common denominaotr, which in this problem, is 72 72x = -64y + 45 -- subtract 45 from both sides 72x - 45 = -64y -- divide both sides by -64 (72x - 45) / -64 = y or... -72/64x + 45/64 = y - 9/8x + 45/64 = y ================= oohhh...sorry...lost track of time...got to go.....I will prob be back on later

OpenStudy (anonymous):

I would almost assume that there is a flaw in the answer they have online, because three people got that same answer. I'm not sure, can your class explain the correct way to do it?

OpenStudy (anonymous):

If i ask it to show me how to do it it gives me a different question, So i keep getting different questions that I dont know how to answer 0_0 @ninjadevo @texaschic101